Find the product of eigen value of matrix

1
Expert's answer
2022-02-07T14:18:31-0500

From the property of eigenvalues, the product of eigenvalues of a matrix equals the determinant of the matrix.

Now, the given matrix is;

"\displaystyle
P=\begin{bmatrix}
 2 & 0&1 \\
 4 & -3&3\\
 0&2&-1
\end{bmatrix}", and "\displaystyle
|P|=\begin{vmatrix}
 2 & 0 & 1 \\
 4 & -3 & 3\\
 0 & 2 & -1
\end{vmatrix}=2\begin{vmatrix}
 -3 & 3 \\
 2 & -1
\end{vmatrix}+1\begin{vmatrix}
 4 & -3 \\
 0 & 2
\end{vmatrix}=2(3-6)+1(8-0)\\\quad\ \ =2(-3)+1(8)=-6+8=2"

Thus, the product of the eigenvalues of "\displaystyle
P=\begin{bmatrix}
 2 & 0&1 \\
 4 & -3&3\\
 0&2&-1
\end{bmatrix}"is "2."